40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 263 of 354  Not logged in ELOG logo
ID Date Author Type Category Subject
  4632   Thu May 5 04:38:20 2011 KojiSummaryLSCComparison between S3399 and FFD-100

Comparison between Hamamatsu S3399 and Perkin Elmer FFD-100

These are the candidates for the BB PD for the green beat detection as well as aLIGO BB PD for 532nm/1064nm.

FFD-100 seems the good candidate.

 

Basic difference between S3399 and FFD-100

- S3399 Si PIN diode: 3mm dia., max bias = 30V, Cd=20pF

- FFD-100 Si PIN diode: 2.5mm dia., max bias = 100V, Cd=7pF

 

The circuit at the page 1 was used for the amplifier.

- FFD-100 showed 5dB (= x1.8) larger responsivity for 1064nm compared with S3399. (Plot not shown. Confirmed on the analyzer.)

- -3dB BW: S3399 180MHz, FFD-100 250MHz for 100V_bias. For 30V bias, they are similar.

Attachment 1: PD_response.pdf
PD_response.pdf PD_response.pdf PD_response.pdf
  4631   Thu May 5 00:08:59 2011 JenneUpdateLSC(Almost) New Screens for RFPDs

I modified C1LSC.mdl to use the CDSphase blocks, which automatically calculate the R and D phase rotation for us.  Now each of the RFPDs has 2 channels in place of the old IQ_MTRX channels:  C1:LSC-RFPD_PHASE_R and C1:LSC-RFPD_PHASE_D.

I have not yet compiled / rebooted / done CDS magic to actually make these installed.  So far the change is only in the simulink model.

I was going to wait until morning to compile/reboot/magic, so I can do it under Joe's supervision.

In the meantime, I also modified the RFPD screens.  They have white boxes for the _R and _D channels just now, but that's because the new model hasn't been put in.  They now look like phase rotators, instead of Koji's temporary matrix.

Still to do:  Find the EPICS database where the phase rotation calculation is done (you give it an angle, it gives you sin(angle) and cos(angle) ).  I want to put a "90-angle" in the database so that we can type in the measured relative phase between I and Q, and it will calculate how many more degrees it needs to get to 90deg. 

 

  4630   Wed May 4 17:32:06 2011 Larisa ThorneConfigurationElectronicsJenne-Seismometer LPF project

Building on what was posted previously

 

 

The configuration has now evolved into an Inverting Op Amp Feedback Low Pass Filter circuit.

Had to change out some components to satisfy conditions: R1=1k Ohm, R2=10k Ohm, C=0.1uF. These were changed in order to decrease the magnitude of the current passing through the op amp by a factor of 10 (10V supplied through the R1 resistor yields about 10mA). The configuration itself was changed from non-inverting to inverting in order to get the frequency vs. gain part of the Bode plot to continue to decrease across higher frequencies instead of leveling off around 4kHz.

Attachment 1: SeisLPF3.jpg
SeisLPF3.jpg
  4629   Wed May 4 15:56:09 2011 valeraSummaryGeneralPSL and MC trends

The attached plot shows 2 day trends of the PMC and MC reflected and transmitted power, the PSL POS/ANG QPD signals, and the temperature measured by the dust counter.

The power step in the middle of the plot corresponds to Koji/Jenne PMC realignment yesterday.

It looks like everything is following the day/night temperature changes.

Attachment 1: pslmcdrift.pdf
pslmcdrift.pdf
  4628   Wed May 4 15:39:32 2011 SureshUpdateRF SystemRF Source Harmonics


I have measured the RF source harmonics in dBm using the HP 8591E spectrum analyser. There is a small discrepancy (< 1 dBm) in the value of RF power shown by the power meter and the HP8591E. This is probably due to the loss of calibration over time.

Initial problem I faced was that when we try to measure the weak harmonics, many below -50dBm we have to choose a small band as advised by Rana. However due to the large amplitide of the fundamental typically around 15dBm or so, the preamp on the spectrum analyser becomes saturated and nonlinear. This gives rise spurious harmonics not present in the source but are rather an aritifact of measurement. The power in harmonics to avoid this I used filters to selectively attenuate the fundamental component (11 or 55 MHz) and then measure the weak harmonics.

However the filters proved difficult to use, because over their stop-band they do not have an input impedance of 50 Ohm. As a result they produce unreliable power measurements for those frequency components which are within the stop band.

To get around this problem I used a suitable attenuator so that even when the internal attenuation is decreased the preamp does not saturate

All the measurements are recorded in the attached document. Pages 4 and 5 give the reliable measurements with the attenuator.

Notes:
1) At times we can see the 29.5 MHz component reflected back from the triple resonant EOM driver.
2) In the 29.5 MHz source output there is a forest of peaks around 100 MHz, which disappear after passing through the AM stabiliser. This suggests that they are associated with AM modulation and have been removed by the stabilizer. But I did not check this further.










Quote:
You should be able tosd resolve the other harmonics by decreasing the IF BW or RBW on the analyzer. Even though
they're OK, its useful to have the final measurement of all of them in some kinds of physical units (like dBm, but
not dBm/Hz or dB or dBcubits).
Attachment 1: RF_Harmonics_Sheet1.pdf
RF_Harmonics_Sheet1.pdf RF_Harmonics_Sheet1.pdf RF_Harmonics_Sheet1.pdf RF_Harmonics_Sheet1.pdf RF_Harmonics_Sheet1.pdf RF_Harmonics_Sheet1.pdf
  4627   Wed May 4 15:22:39 2011 kiwamuUpdate Re: Intermittent MC3 UL PD signal

I went push all the possible connectors for the MC3 shadow sensors including the SCSIs, flat cables and satellite box.

Also I put screws on them so that they won't become loose any more.

As a result UL_PDMON dropped from 0.6 V to 0.490 V and it becomes stable so far.

I didn't strain relief the cables but we must do it at some point before going into the full locking test.

Quote from #4625

The attached plot shows the 30 day trend of the MC3 UL PD signal. The signal dropped to zero at some point but now it is close to the level it was a few weeks ago. There still could be a problem with the cable.

The rest of the MC1,2,3 PD signals looked ok.

 

  4626   Wed May 4 13:57:04 2011 kiwamuUpdateSUSmechanical resonances updated

[Leo  w/ a little help from Kiwamu]

Leo summarized the mechanical resonances of all the suspensions, based on the free-swinging spectra taken on Sat Apr 30.

Since Leo doesn't have the wiki account I helped him putting the information on the wiki.

Good work, Leo !

http://blue.ligo-wa.caltech.edu:8000/40m/Mechanical_Resonances

Quote from

Here are the free-swinging spectra for the BS, ETMX, ETMY, ITMX, ITMY, MC1, MC2, MC3, and PRM chambers.  Kiwamu left the suspensions free for 5 hours this weekend, starting at Sat Apr 30 00:15:26 2011.

 

  4625   Wed May 4 13:51:51 2011 valeraConfiguration Intermittent MC3 UL PD signal

The attached plot shows the 30 day trend of the MC3 UL PD signal. The signal dropped to zero at some point but now it is close to the level it was a few weeks ago. There still could be a problem with the cable.

The rest of the MC1,2,3 PD signals looked ok.

Attachment 1: mc3ulpdmon.pdf
mc3ulpdmon.pdf
  4623   Wed May 4 13:45:56 2011 kiwamuUpdateLSClocking last night

Last night I was trying to calibrate the MICH error signal and the actuators on BS and ITMs.

However I gave up taking the data because the MC locking was unstable. MC3 drifted a lot.

  4622   Wed May 4 12:07:48 2011 SureshUpdateRF SystemREFL55 installed on the AP table

REFL55 has been installed on the AP table.  REFL11 has been moved to make space for a 50% beam splitter. The reflected beam from this splitter is about 30% of the transmitted beam power.  The reflected beam goes to REFL11 in the current configuration.  The DC levels are 1.2V on REFL 11 and 3.5V on the REFL55.

I redid some of the cabling on the table because the we need to choose the heliax cables such that they end up close to the demod board location.  As per the 1Y2 (LSC) rack layout given here, some of the PD signals have to arrive at the top and others at the bottom of the LSC rack.

Currently the PDs are connected as follows:

 

REFL11 PD --> Heliax (ASDD133) (arriving at the top of LSC rack) --> REFL11 Demod Board 

REFL55 PD --> Heliax (REFL166) (arriving at the top of LSC rack) --> AS55 Demod Board

AS55 PD --> Heliax (AS166) (arriving at the top of the LSC rack) --> not connected.

 

We are waiting for the Minicircuits parts to modify the rest of the demod boards.

 

The heliax cables arriving at the LSC rack are not yet fixed properly.  I hope to get this done with Steve's help today.

 

 

  4621   Wed May 4 11:48:01 2011 SureshUpdateIOOMC Locking not working

[Valera, Suresh]

The first time I noticed that the MC was not locking was after I had finished switching the RF source installation.  Before this change the RF modulation frequency (for MC) was 29.485 MHz as read from the Marconi RF Source.  We replaced this with a Wenzel crystal source at 29.491 MHz.  This may have changed the loop gain. 

Today, I changed the MC alignment to optimise the MC lock.  Valera pointed out that this is not a desirable solution since it would shift the beam pointing for all components downstream.  However, since we are not sure what was the last stable configuration, we decided to stay with the current settings for now and see the trends of several parameters which would tell us if something is drifting and causing the autolocker to fail.

The MC Auto locker is now working okay.  However to obtain lock initially we reduced the loop gain by decreasing the VCO gain.  We then increased the gain after the autolocker had locked the MC.

 

 

 

  4620   Tue May 3 18:46:06 2011 ranaUpdateIOOMC Locking not working

I found that the MC autolocker was OFF. Kiwamu says he turned it off because its slow. Suresh says that he has some feelings that maybe something is wrong. I'll let them describe what they know about the MC in an elog.

I checked the trend of the MC and PMC transmissions for the past 30 days:

Untitled.png

Looks like the alignment has been drifitng. PMC was corrected recently by Koji, but the alignment of the input beam to the MC or the MC itself has to be fixed. Has someone been twiddling the MC SUS alignment biases??

Attachment 1: Untitled.png
Untitled.png
  4619   Tue May 3 18:25:38 2011 kiwamuSummaryLSCPRMI locking : plan

[Rana/Jamie/Kiwamu]

 Since we've got the PRMI locked we now should be able to do more qualitative measurements.

Here is a task list that we will measure/develop in the PRMI condition.

 

 

 - Optical gain measurements

 - Characterization of control loops

 - MICH and PRC calibrations

 - Noise budget

 - Development of automatic noise budget scripts

 - Arm loss measurement 

 - Shnupp asymmetry measurement

  4618   Tue May 3 17:19:25 2011 kiwamuUpdateElectronicsHeliax connectors on 1Y2 rack : tightened

My observation wasn't accurate enough.

The looseness came from the fact that the N-SMA bulk heads were slipping on the black plate.

This is actually what Suresh pointed out (see here). So the thickness of the black plate doesn't matter in this case.

Somehow I was able to tighten the bulk heads using two wrenches and I think they are now tight enough so that the heliax's heads don't move any more.

Quote from #4601

I found that all the Heliax cables landing on the bottom of 1Y2 were too loose.

The looseness basically comes from the fact the black plate is too thick for the Heliax cable to go all the way. It permits the Heliax's heads to rotate freely.

 

  4617   Tue May 3 16:47:55 2011 kiwamuUpdateLSCPRMI locking : details

Here are some details about the PRMI locking done last night.

 

(REFL11 installation)

 REFL11 has been installed on the AP table. The RF signal from the RFPD is sent by a heliax cable which has been called ASDD133.

Before the beam goes into the RFPD a HWP and PBS are installed such that we can adjust the amount of light entering to the photo diode.

One thing I didn't like was that I had to introduce a big amount of the light into the PD to get a reasonably big RF signal.

I was trying to look for an RF signal by looking at a spectrum analyzer, then I realized that the RF signal at 11 MHz was quite tiny when the DC_MON was less than 1.5 V.

After I increased the amount of the light up to 1.9 V in DC_MON, which sounds already too much, I then got able to see the 11 MHz signal on the analyzer.

Note that I decreased the amount of the light down to 0.5 V after I finished locking the PRMI.

We should make sure what is going on with the 11 MHz modulation.

APtable.png

 

(Locking)

First I started locking the MIchelson with AS55. The demodulation phase was already somewhat optimized to the I-signal port.

So I decided not to touch the demodulation phase matrix because it may take some times.

After I eliminated electrical offsets in the digital side, I was easily able to lock the Michelson. The control sign was plus.

Then I started playing with the PRC control too. The demodulation phase in REFL11 looked nearly 45 deg although I didn't carefully measure it.

I made a 45 deg rotational matrix to maximize the I-port signal and tried to lock the PRC. Then immediately I was able to lock PRC as well as MICH.

 GAIN_MICH = 100

 GAIN_PRC = 100

Also GAIN_PRC = -100 gave a carrier resonant lock.

The control filters are the same in MICH and PRC. I used my favorite filters as usual.

 FM1 = 1000 : 10

 FM6 = 0.1 : 1

 FM7 = 1 : 50

Somehow I frequently failed to engage the boost filters (i.e. FM6 and FM7) it looks offsets in the control path kicks either BS or PRM.

Quote from #4604

The PRMI has been successfully locked

Details will be posted in the morning. 

 

  4616   Tue May 3 16:20:13 2011 steveUpdateSUSPRM & BS oplevs are ready

Quote:

The returning spot diameter on the qpd ~10 mm. In order to reduce the spot size I moved the f 1145 mm lens toward the PRM ~ 25 cm. The spot size was reduced to ~8 mm, 3200 counts.

I'll try to find an other lens tomorrow.

 Atm 1,  PRM oplev inward path with 2 lens solution: 14 cm gap between F 1145 and F 1545 mm lenses. 

Atm 2,   The PRM beam size 3 mm and the beam  quality is still bad. The BS path only needed alignment.

Attachment 1: P1070630.JPG
P1070630.JPG
Attachment 2: P1070632.JPG
P1070632.JPG
  4615   Tue May 3 15:59:22 2011 steveFrogsPhotosX-mas comes early

The little red all terrain cargo wagon 40" x 18"  has just arrived on pneumatic wheels.

Model #29, 200 lbs max load at 26 PSI,  minimum age requirement 1.5 years

Attachment 1: P1070634.JPG
P1070634.JPG
  4614   Tue May 3 15:48:26 2011 Larisa ThorneUpdateElectronicslaser temperature control LPF, final version!

This is a continuation of this

 The low pass filter is finally acceptable, and its Bode graph is below (on a ~3Hz frequency span that shows the cutoff frequency is at 0.1Hz)

Attachment 1: LPF100k.2.jpg
LPF100k.2.jpg
  4613   Tue May 3 15:04:20 2011 kiwamuUpdateSUSRe: DRMI prep : suspension diagnostic

Jenne went through all the suspension racks and pushed all the connectors.

After pushing them, we had a quick look at those spectra and found no funny noise spectrum except for C1:PRM-SENSOR_UL.

We then checked connection around the SCSI cables and eventually found the connection between ADC_card_0 and a SCSI was loose.

We put short standoffs on the ADC card so that the screws from the SCSI can nicely reach to the ADC card. Now everything looks fine.

SUS diagnostic is quite useful !

Quote from #4611

Notice that the C1:SUS-ITMX_SENSOR_UL and C1:SUS-MC3_SENSOR_UL spectra fall as 1/f.  Jenne suggested that this might indicate that there is a loose electrical connection.

  4612   Tue May 3 14:35:44 2011 Larisa ThorneConfigurationElectronicsJenne-Seismometer LPF project

 I was charge with making a Non-Inverting Op Amp Low Pass Feedback circuit for Jenne, which may somehow be integrated into a seismometer project she's working on.

 

Circuit diagram is attached. Calculations show that R1, R2 and C have the following relationship: if R1=10^n, R2=10^(n+1), C=10^(-n-4). For the particular circuit being modeled by the transfer function, R1=100 Ohm, R2=1k Ohm, and C=1uF.

Attached also is the circuit's Bode plot, showing frequency versus gain and phase, respectively. The frequency versus gain graph is true to what the circuit was calculated to generate: a gain of +20 and a cutoff frequency at 200Hz. Not sure what's going on with the frequency verus phase plot.

Attachment 1: SeisLPFdiagram.jpg
SeisLPFdiagram.jpg
Attachment 2: seisLPF.Bose.jpg
seisLPF.Bose.jpg
  4611   Tue May 3 13:22:13 2011 LeoUpdateSUSRe: DRMI prep : suspension diagnostic

Here are the free-swinging spectra for the BS, ETMX, ETMY, ITMX, ITMY, MC1, MC2, MC3, and PRM chambers.  Kiwamu left the suspensions free for 5 hours this weekend, starting at Sat Apr 30 00:15:26 2011.

This is GPS time 988 182 941.  Quick tip: you can do local to GPS time conversions using lalapps_tconvert, which is a lot like tconvert but with special powers.  It is installed on pianosa.

$ lalapps_tconvert Sat Apr 30 00:15:26 2011

988182941

I generated these figures with the attached Python script, measure.py.

Notice that the C1:SUS-ITMX_SENSOR_UL and C1:SUS-MC3_SENSOR_UL spectra fall as 1/f.  Jenne suggested that this might indicate that there is a loose electrical connection.

Also, notice that C1:SUS-ETMY_SENSOR_LR, C1:SUS-ITMY_SENSOR_LL, and C1:SUS-PRM_SENSOR_SIDE are a lot noisier above 10 Hz.

Attachment 1: BS.png
BS.png
Attachment 2: ETMX.png
ETMX.png
Attachment 3: ETMY.png
ETMY.png
Attachment 4: ITMX.png
ITMX.png
Attachment 5: ITMY.png
ITMY.png
Attachment 6: MC1.png
MC1.png
Attachment 7: MC2.png
MC2.png
Attachment 8: MC3.png
MC3.png
Attachment 9: PRM.png
PRM.png
Attachment 10: SRM.png
SRM.png
  4610   Tue May 3 11:49:03 2011 KojiUpdateLSCdaytime tasks

Done. C1:PSL-PMC_PMCTRANSPD was improved from ~0.75 to 0.87.

Quote:

- PMC alignment (Jenne/Koji)

 

  4609   Tue May 3 10:59:31 2011 josephbUpdateCDS1Y2 binary output adapter board now powered

I temporarily turned off the power to the 1Y2 rack this morning while wiring in the binary output adapter board power (+/- 15V) into the cross connects.

The board is now powered and we can proceed to testing if can actually control the LSC whitening filters.

  4608   Tue May 3 10:41:35 2011 josephbUpdateCDSMorning maintenance

1) Filled in the C1SUS_BS_OLMATRIX properly so as to make the BS oplev work for Steve.

2) Turned on the ITMX damping.  Apparently it had tripped this morning, possibly due to work in the lab area.

3) The ETMX FE controller (c1scx) had ADC timed out and died sometime around 8:30 am.  The c1x01 (the IOP on the ETMX computer) was also indicating a FB status error (mismatch in DAQ channels).

The reported error in dmesg on c1iscex was:

[1628690.250002] c1spx: ADC TIMEOUT 0 3541 21 3605
[1628690.250002] c1scx: ADC TIMEOUT 0 3541 21 3605

Just to be safe, I rebuilt the c1x01 and c1scx models, ran ./activateDAQ.py, and used the scripts killc1spx, killc1scx, and killc1x01.

I finally restarted the process with startc1x01, startc1scx, and startc1spx.  Everything is currently alive and indicating all green.

  4607   Tue May 3 10:21:25 2011 KojiUpdateLSCdaytime tasks

I think the installation of the PD DC signals are quite important. What to do
1) Connect the DC signals to the right top whitening board (be aware that there may be the modification of the whitening circuit).
2) Reconfigure the LSC model such that the DC signal is passed to the right channels (modify the left top part of the model)

Quote:

Daytime tasks :

 - PRM & BS oplev (Steve)

 - LSC binary outputs (Joe/Jamie)

 - installation of the REFL55 RFPD (Suresh/Jamie)

 - Adjustment of demodulation phases (Kiwamu)

 - Bounce-Roll filters on BS and PRM (Suresh/Joe)

 - Suspension diagnostic using the free-swinging spectra (Leo)

 - PMC alignment (Jenne/Koji)

 

  4606   Tue May 3 05:32:04 2011 kiwamuUpdateLSCdaytime tasks

Daytime tasks :

 - PRM & BS oplev (Steve)

 - LSC binary outputs (Joe/Jamie)

 - installation of the REFL55 RFPD (Suresh/Jamie)

 - Adjustment of demodulation phases (Kiwamu)

 - Bounce-Roll filters on BS and PRM (Suresh/Joe)

 - Suspension diagnostic using the free-swinging spectra (Leo)

 - PMC alignment (Jenne/Koji)

  4605   Tue May 3 04:09:53 2011 KojiUpdateLSCREFL55 PD update

[Rana, Koji]

REFL55 was modified. The noise level confirmed. The PD is now ready to be installed.

 


Kevin's measurement report told us that something was wrong with REFL55 PD. The transimpedance looked OK, but the noise level was terrible (equivalent to the shotnoise of 14mA DC current).

Rana and I looked at the circuit, and cleaned up the circuit, by removing unnecessary 11MHz notch, 1k shunt resister, and so on.

I made a quick characterization of the PD.

First page:

The transimpedance ws measured as a function of the frequency. The resonance was tuned at 55MHz. The notch was tuned at 110MHz in order to reject the second harmonics. The transimpedance was ~540V/A at 55MHz. (For the calibration, I believed the DC transimpedance of 50V/A and 10000V/A for the DC paths of this PD and #1611, respectively, as well as the RF impedance (700V/A0 of #1611.

Second page:

Output noise levels were measured with various amount of photocurrent using white light from a light bulb. The measurement was perforemed well above the noise level of the measurement instruments.

Third page:

The measured output noise levels were converted into the equivalent current noise on the PD. The dark noise level agrees with the shot noise level of 1.5mA (i.e. 22pA/rtHz). In deed, the noise level went up x~1.5 when the photocurrent is ~1.4mA.

Attachment 1: REFL55.pdf
REFL55.pdf REFL55.pdf REFL55.pdf
  4604   Tue May 3 03:53:20 2011 kiwamuUpdateLSCPRMI locking : done

The PRMI has been successfully locked

PRMI_locking.png

Also changing the sign of the PRC control gave me the lock of the carrier resonant condition.

The screenshot above is the time series of the error signals when I was locking the PRMI in the sideband resonant condition (i.e. carrier is non-resonant).

Note that I used REFL11 for the PRC control and AS55 for the MICH control as planed.

Details will be posted in the morning.

Quote from #4595

Today we will try to lock the PRMI.

 

  4603   Tue May 3 00:44:02 2011 KojiConfigurationComputersMartian WIreless Bridge

The Martian wireless bridge has the ethernet cable inserted in the wrong connector.

It should be inserted to one of the four port. Not in the "INTERNET" connector.

Once the connector has been changed, the martian net as well as the internet became accessible from the laptops.

  4602   Mon May 2 23:29:57 2011 SureshUpdateElectronicstoo loose Heliax connectors on 1Y2 rack

The countersink gives rise to another problem when we mount the N-type-to-SMA bulkhead adaptor.  As we are making a circular hole in the plastic strip (instead of a hole with two flat sections)  the adaptor is free to turn when we tighten it with a wrench.  We currently hold the smooth circular part on the other side with a gripping pliers and while tightening.  If that part disappears into the countersink (as seen in the pics) we will not be able to tighten the adaptor sufficiently and consequently we will also not be able to get the heliax connector to be tight.

A better solution would be to use the 1/4-inch plastic L-angle beam which Steve has used on the AS table.  In addition to solving this loose connector problem, the beam is also more rigid than the plastic strip.

 

  4601   Mon May 2 21:14:55 2011 kiwamuUpdateElectronicstoo loose Heliax connectors on 1Y2 rack

I found that all the Heliax cables landing on the bottom of 1Y2 were too loose.

Due to this loose connection the RF power at 55 MHz varies from -34 dBm to 3 dBm, depending on the angle of the Heliax's head.

DSC_2982_ss.jpg

 

The looseness basically comes from the fact the black plate is too thick for the Heliax cable to go all the way. It permits the Heliax's heads to rotate freely.

What we should do is to make countersinks on the black plate like this:

DSC_2984_ss.jpg

 

  4600   Mon May 2 18:16:24 2011 kiwamuUpdateLSCPRMI locking : status update

(done)

- AS55 RFPD

  With a help from Jamie the AS55 RFPD was installed.

- 55 MHz demodulation board

  The AS55 demod board was installed on 1Y2.

- 3-way combiner

  ZFSC-3-13 has been installed. All the RF cables from the source side were connected to the combiner.

 

(next things)

 - installation of the REFL11 RFPD

 - DAQ check for AS55 and REFL11

Quote from #4595

(to be done in the daytime)

 - setup AS55 RFPD

 - install a demod board for 55 MHz

 - install a 3-way RF combiner on EOM.

 - prepare 55 MHz RF source (Marconi or RF source box ?)

 

  4599   Mon May 2 17:10:55 2011 steveUpdateSUSPRM oplev

The returning spot diameter on the qpd ~10 mm. In order to reduce the spot size I moved the f 1145 mm lens toward the PRM ~ 25 cm. The spot size was reduced to ~8 mm, 3200 counts.

I'll try to find an other lens tomorrow.

  4598   Mon May 2 15:53:27 2011 kiwamuUpdateLSCPRMI locking prep : demodulation boards

I swapped the name of two demodulation boards (AS55 and REFL55).

Now the REFL11 and AS55 demodulation boards are ready to go for the PRMI locking.

The physical labels, which are on the front surface of the boards, are also corrected to avoid a confusion.

Here is the latest RF status.

RF_Work_Status.png

Quote from #4574

The files are on https://nodus.ligo.caltech.edu:30889/svn/trunk/suresh/40m_RF_upgrade/.

 

  4597   Mon May 2 13:43:05 2011 steveFrogsPhotosbirthday boys

.....Happy.... Birthday.... to.... Joseph... and... Jamie...Happy....Birthday..... to.... You............sing with us........Happy Birthday.....to you

Attachment 1: P1070622.JPG
P1070622.JPG
  4596   Mon May 2 13:30:13 2011 kiwamuUpdateSUSPRMI locking prep: PRM oplev

I had a quick look at PRM optical lever.

The He-Ne beam is still successfully coming out from the chamber and I could guide it to the QPD by using steering mirrors.

But the beam size looks too big for the QPD. We should slide the lens which is standing before the injection to get a moderately smaller beam size at the QPD.

Quote from #4595

 - activation of PRM oplev

  4595   Mon May 2 12:51:52 2011 kiwamuUpdateLSCPRMI locking : plan

Today we will try to lock the PRMI. Here is a plan for it.

(to be done in the daytime)

 - setup REFL11 RFPD

 - setup AS55 RFPD

 - install a demod board for 55 MHz

 - install a 3-way RF combiner on EOM.

 - prepare 55 MHz RF source (Marconi or RF source box ?)

 - adjustment of each demodulation phase

 - activation of PRM oplev

 

 (control topology)

- AS55_Q ==> BS

- REFL11_I ==> PRM

  4594   Mon May 2 11:14:27 2011 kiwamuUpdateSUSscript : opticshutdown

Just FYI.

There is a useful script for this particular job : shutting down all the suspensions and bringing it back to operation after 5 hrs.

It is called opticshudown, which resides in /cvs/cds/rtcds/caltech/c1/scripts/SUS/.

 

Also I added this script on the list in the wiki where all the scripts will be listed.

http://blue.ligo-wa.caltech.edu:8000/40m/Computers_and_Scripts/All_Scripts#opticshutdown

If you find any other useful scripts, please add them on the wiki.

Quote from #4592

I leave all the suspensions free from the watchdogs for 5 hours from now.

 

  4593   Sat Apr 30 05:14:33 2011 KojiUpdateGreen LockingElectronics schematic for vertex beatbox

- AC coupling for the comparator circuit of the green locking

In order to relieve the power consumption of the RF buffer, ac coupling circuits have been added.

The ac coupling before the buffer amp helps to relieve the power consumption in the chip.
But because of the distortion of the signal (and the limitation of the bandwidth), the output still has some DC (~0.6V).
Therefore, the output is also AC coupled.

Note that the BW pin of BUF634P should be directly connected to -15V in order to keep the bandwidth of the buffer.

The drawings are also uploaded on the green electronics wiki

Attachment 1: freq_divider.png
freq_divider.png
  4592   Sat Apr 30 00:15:26 2011 kiwamuUpdateSUSDRMI prep : suspension diagnostic

As a part of the DRMI preparation,

I leave all the suspensions free from the watchdogs for 5 hours from now.

Please DO NOT touch them.

I will check the spectra and the mechanical resonant frequencies on Monday.

Also I will renew all the input matrices of the local dampings based on these free swinging spectra.

  4591   Fri Apr 29 18:24:05 2011 SureshUpdateRF SystemRF system: 1X2 Rack cabling

[Joe, Jamie, Suresh]

We have installed the IDE to SCSI adaptor module into the 1X2 rack and have connected the AA filter outputs to it.

P4290070.JPG

 

We have removed the following cables running between the 1X2 and 1X3 racks.

The long twisted pair ribbon cable which previously carried the ADC signals.

1X2-ASC 6, 1X2-ASC 47, 1X2-ASC 9, 1X2-ASC 8, 1X2-ASC 10, 1X2-ASC 7,

CAB-1X2-LSC 42, CAB 1X2-LSC 56,  CAB 1X2-LSC 41, CAB 1X2-LSC 43

1X3-2 ASC 47

We have also removed the following by mistake.  We will put them back them on Monday

1X2-LSC 21, 1X2-LSC-20.

We have also removed the ASC QPD cables and moved the QPD cards which were present in the middle Eurocate (#2) to the unused Eurocrate at the bottom position (#3).

The binary input cables at the back of the cards require to be supported so that their weight does not pull them out of the sockets at the back of the crates.

Some of the slots where we plan to plug in Demod boards (the 165 MHz boards)  are not currently connected to any binary output on the C1:LSC computer.  We need these binary controls for the fitlter modules on the cards.

When we eventually begin to use the 15PDs as planned, then we will occupy 30 ADC channels (I & Q outputs).  Currently we have just one ADC card installed on the C1:LSC providing 32 ADC channels.  Joe found another 16bit 32 channel ADC card in his stash but we need to get a timing+adaptor board for it. In general we are going to need the third Eurocrate.

A platform for the power supply of the RF Distribution box needs to be built and the power supply needs to be moved into the 1X2 rack rather than sit on top of 1X2 rack.

 

 

 

  4590   Fri Apr 29 14:36:36 2011 josephb, steveUpdateCDSRemoved hanging D-sub to SCSI in 1X5

Quote:

Problem:

Way back, Jay had D-sub to SCSI adapters made to adapt our existing Sander box AA filters to the new SCSI based IO chassis.  However, these did not fit inside the box.

At the time, we simply left the cards outside hanging, which was a hack and needed to be replaced.

Solution:

Steve modified a black AA filter box so that it could fit the D-sub to SCSI adapter board on it, plus strain relief the SCSI cable, rather than let it hang.  The back of the box was cut, and an extending piece of metal attached to the bottom of the box.  The adapter board was screwed into the box, the SCSI plugged in, then the SCSI cable is clamped to the extending metal as well.

This modification will be propagated to the 3 remaining AA filter boards using the D-sub to SCSI adapter.

 The same modification was carried out at 1X5 for PRM & SRM.

Note:  D68L8EX-850Hz  are removed  and bypassed in 7 channels.

Attachment 1: P1070621.JPG
P1070621.JPG
  4589   Fri Apr 29 13:58:36 2011 ranaUpdateLSCY arm locked : details

Quote:

 * Farfalla, a lab laptop, seems out of network.

 If you look at the real host table instead of the misleading host table in the wiki, you will see that someone has deleted Farfalla from there. She needs to be re-added.

  4588   Fri Apr 29 13:15:04 2011 kiwamuUpdateLSCY arm locked : details

As far as I know, this button works only once after the launch of MEDM...

Quote:

 * reload button on sitemap.adl doesn't work. 

 

  4587   Fri Apr 29 12:18:48 2011 kiwamuUpdateLSCY arm locked : details

First of all, the conclusions / results from the exercise of the Y arm locking yesterday are:

   The position of the beam spots on both ETMY and ITMY are now not so bad ( ~ 5 mm off from the center).

   The input PZTs are coarsely aligned to the Y arm.

   Nevertheless IP_ANG is still too high to come out from the view port at the Y end station.

  After the alignments of PRM, SRM and Michelson, POP is still largely clipped.

 


(what I did)

  - Alignments of the Input PZTs

    First I tried letting the incident beam hit the center of ETMY by steering PZT1 and 2 as usual.
    Then I coarsely aligned the cavities to the incident beam and checked the beam flashing spots on ETMY and ITMY with the CCD monitors.
    When the spots were far from the center I went back to the alignment of PZT1 and 2 to get better beam positions. And repeated this work several times.

  - Adjustment of the demodulation phase  for the Y arm PDH.

    First I started looking at the digital signals and tried correcting the demodulation phase by the rotation matrix, but this didn't go fast because I had to do some DAQ settings, plotting and analysis.
    Instead looking at the digital signal, I observed the analog signals with an oscilloscope. I found the demodulation phase was something like 45 deg.
    Based on the analog measurement I rotated the digital matrix by 45 deg to get the I-signal maximized. Indeed this worked well. I obtained a beautiful PDH signal from the I-signal.
    Note that we are using 11 MHz mod/demod and eventually the signals come out from "REFL33" on the digital side.

  - Activation of oplev on ITMY

    Instead using ETMY I used ITMY for the length control because somehow I felt that ETMY coils were suspicious and they looked not so nice.
    One of the reason is that ETMY's coil actuation efficiencies looked low compared to the other test masses.
    For example a gain of 700 for SUSSIDE damping is needed on ETMY to get a reasonable Q. This is about 2 - 3 times larger gain than the other test masses.
    So I started using ITMY for the locking and activated the oplev to suppress unwanted excitation due to kicks from the control signal during the locking,
    The oplev has been misaligned, so I went to the ITMY optical bench and tweaked a steering mirror to let the He-Ne beam go into the QPD.
    I set the gains +2 for PITCH and -2 for YAW.

  - PDH locking

    The locking had been quite difficult even though the cavity alignment was quite good.
    It's because the beam on AS11_RFPD was almost falling off from the photo diode. This causes a big amplitude fluctuation in the PDH signal as the beam position moves.
    After aligning the beam by steering BS I got able to lock the Y arm. The PD whitening gains are all 0 dB and the feedback gain is -2, giving us a UGF of 250 Hz.

 

(Broken or likely broken stuff)

 * IP_ANG doesn't give a signal to the digital side.

 * ETMY coils look weak and 2 - 3 times weaker than the other test masses. (or OSEM readout gain maybe lower)

 * reload button on sitemap.adl doesn't work.

 * Farfalla, a lab laptop, seems out of network.

Quote from #4586

The Y arm has been locked with the IR beam. The purpose is to use the arm as an alignment reference for the input PZTs.

Detail will be posted later. Here is a picture of ITMY suspension. You can see there is a beam spot in the middle of the test mass.

 

  4586   Fri Apr 29 05:48:52 2011 kiwamuUpdateLSCY arm locked

The Y arm has been locked with the IR beam. The purpose is to use the arm as an alignment reference for the input PZTs.

Detail will be posted later. Here is a picture of ITMY suspension. You can see there is a beam spot in the middle of the test mass.

DSC_2978_ss.jpg

  4585   Fri Apr 29 03:39:49 2011 KojiSummaryLSCCavity lengths

I tried the idea that the PRC can resonate f1 and f2 at the same time if the arm gives the reflection phase to f1 and f2 with the ratio of 1 vs 5.

The details are described on wiki. The point is this removes all of the PRC/SRC/asymmetry mumbo jumbo.

The calculated cavity lengths for f_mod of 11.065399MHz are:

  • Arm Length: 37.7974 [m]

  • PRC Length: 6.7538 [m]

  • SRC Length: 5.39915 [m]

  • Asymmetry (lx-ly): 0.0342 [m]


Here is the actual values derived from the photos.

  • Arm Length: 37.54 [m] (0.26m too short)

  • PRC Length: 6.760 [m] (6mm too long)

  • SRC Length: 5.415 [m] (16mm too long)

  • Asymmetry (lx-ly): 0.0266 [m] (8mm too long)

  4584   Thu Apr 28 22:38:38 2011 AidanUpdateGreen LockingElectronics schematic for vertex beatbox

 With some assistance from Kiwamu and Koji, I've drawn up the electronics design for the Beat Box for the vertex green locking. The Omingraffle schematic is posted on the Green Locking Wiki page. It's also attached below. Some final touches are necessary before we can Altium this up.

 Attachment 1: Schematic of beatbox

Attachment 2: Front and back panel designs.

Attachment 1: Canvas_1.png
Canvas_1.png
Attachment 2: Canvas_2.png
Canvas_2.png
  4583   Thu Apr 28 16:12:19 2011 josephb, jamieUpdateCDSNew CDS model SVN

New SVN

We are now using the LIGO CDS SVN for storing our control models.

The SVN is at:

https://redoubt.ligo-wa.caltech.edu/websvn/

The models are under cds_user_apps, then trunk, then approriate subsystem (ISC for c1lsc for example), c1 (for caltech 40m), then models.

We have checked out /cds_user_apps to /opt/rtcds/.

So to find the c1lsc.mdl model, you would go to /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1lsc.mdl

This SVN is shared by many people LIGO, so please follow good SVN practice.  Remember to update models ("svn update") before doing commits.  Also, after making changes please do an update to the SVN so we have a record of the changes.

New Practices

We are creating soft links in the /opt/rtcds/caltech/c1/core/advLigoRTS/src/epics/simLink/ to the models that you need to build.  So if you want to add a new model, please add it to the cds_users_apps SVN in the correct place and create a soft link to the simLink directory.

lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1sus.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1sus.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1sup.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1sup.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1spy.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1spy.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1spx.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1spx.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1scy.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1scy.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1scx.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1scx.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1mcs.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1mcs.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1x05.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1x05.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1x04.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1x04.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1x03.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1x03.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1x02.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1x02.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1x01.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1x01.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1rfm.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1rfm.mdl
lrwxrwxrwx 1 controls controls      55 Apr 28 14:41 c1dafi.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1dafi.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1pem.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1pem.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1mcp.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1mcp.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1lsp.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1lsp.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1lsc.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1lsc.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1ioo.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1ioo.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1gpv.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1gpv.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1gfd.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1gfd.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1gcv.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1gcv.mdl
lrwxrwxrwx 1 controls controls      54 Apr 28 14:41 c1ass.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1ass.mdl

  4582   Thu Apr 28 15:31:36 2011 kiwamuUpdateIOOMC PDH lock : readjustment of demodulation phase

Since Suresh has installed the RF source box and changed the cable configuration somewhat,

the demodulation phase for the MC locking became off by about 10 degree.

I changed the length of some cables and obtained a good demodulation phase by the same technique as Suresh and Koji did before (see here for detail).

I maximized the Q signal. The lock of the MC looks healthy.

DSC_2975_ss.jpg

Quote from #4578

RF Source box has been mounted in the 1X2 rack. 

ELOG V3.1.3-